Inventory Analysis A company reports the following: Cost of goods sold Average inventory Determine (a) the inventory turnover and (b) the number of days sales in inventory. Round interim calculations to the nearest dollar and final answers to one decimal place. Assume 365 days a year $224,840 56,210 a. Inventory turnover days b. Number of days sales in inventory
Show Me How Calculator Long-Term Solvency Analysis The following information was taken from Acme Companys balance sheet: Fixed assets (net) Long-term liabilities Total liabilities Total stockholders equity Determine the companys (a) ratio of fixed assets to long-term liabilities and (b) ratio of liabilities to stockholders equity. If required, round your answers to one decimal place. $1,167,400 49,000 1,414,350 1,571,500 a. Ratio of fixed assets to long-term liabilities b. Ratio of liabilities to stockholders equity
media%2Fde7%2Fde712524-2f17-430f-95b9-c4
Show Me How Calculator Earnings per Share and Price-Earnings Ratio A company reports the following: Net income Preferred dividends Shares of common stock outstanding Market price per share of common stock $572,000 $32,000 90,000 $47.4 mine the companys earnings per share on common stock. Round your answer to the nearest cent. Use the rounded answer of requirement a for subsequent requirement, if required. b. Determine the companys price-earnings ratio. Round to one decimal place.
Answer #1

Answer(1): (a): Inventory turnover = Cost of goods sold / Average inventory

Inventory turnover: 224840 / 56210 = 4 times

(b): Number of days sales in inventory = No. of days in a year / Inventory turnover ratio

Number of days sales in inventory: 365 / 4 = 91.25 days

Answer(2):

(a): Ratio of fixed asset to long term liabilities: 1167400 / 449000 = 2.6

(b): Ratio of liabilities to stockholder's equity: 1414350 / 1571500 = .9

Answer(3):

(a): Return on stockholder's equity = (Net Income / Average stockholder's equity) * 100

Return on stockholder's equity: 315000 / 2916667 = 10.799 or 10.8%

(b): Return on common stockholder's equity = (Net Income - Preferred dividend) / Average common stockholder's equity

ROCE: (315000-12600) / 1482353 = 20.3999 or 20.4%

Answer(4):

(a): EPS on common equity = (Net Income - Preferred dividend) / Shares of common stock outstanding

EPS on common equity: (572000-32000) / 90000 = 6

(b): P/E ratio = Market price per share / EPS

P/E Ratio: 47.4 / 6 = 7.9

(Please rate)